### v1

# Modular Periodicity of Random Initialized Recurrences

Abstract
Classical studies of the Fibonacci sequence focus on its periodicity modulo $m$ (the Pisano periods) with canonical initialization. We investigate instead the complete periodic structure arising from all $m^{2}$ possible initializations in $\mathbb{Z}/m\mathbb{Z}$. We discover perfect chiral symmetry between the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$, and observe fractal self-similarity in the extension from prime to prime power moduli. Additionally, we classify prime moduli due to their periodicity, and show that they exhibit weight preservation under modular extension. Furthermore, we define a minima distribution $P(n)$ governed by Lucas ratios, which satisfies the symmetric relation P(n)=P(1n)P(n)=P(1-n). For cyclotomic recurrences, we conjecture explicit counting functions for the number of distinct periods, with connections to necklace enumeration. These findings suggest potential links to Viswanath’s random recurrence, modular forms and L-functions.
Keywords— Fibonacci sequence, Pisano period, cyclotomic polynomial, chiral symmetry, recurrence relation, modular arithmetic
1 Introduction
Linear homogeneous recurrence relations play a fundamental role in number theory and combinatorics. A recurrence of order $k$ is defined as
$a_{n}=r_{1}a_{n-1}+r_{2}a_{n-2}+...+r_{k-1}a_{n-k+1}+r_{k}a_{n-k}=\sum\limits_{i=1}^{k}r_{i}a_{n-i}$ (1)
where the sequence is fully determined by $k$ initial values ($a_{0},a_{1},...,a_{k-1}$) and its characteristic polynomial
$f(x)=x^{k}-r_{1}x^{k-1}-...-r_{k-1}x-r_{k}$ (2)
as shown in [1]. The most celebrated example is the Fibonacci sequence $\{0,1,1,2,3,5,8,13,\ldots\}$, generated by $x^{2}-x-1$ with initial values $a_{0}=0,a_{1}=1$. This sequence grows exponentially, but when reduced modulo $m$, it becomes periodic with period length $\pi(m)$, known as the Pisano period [2]. Determining $\pi(m)$ explicitly remains an open problem in number theory.
However, the classical Pisano problem addresses only a single sequence with fixed initialization. In $\mathbb{Z}/m\mathbb{Z}$, there are $m^{2}$ possible initial conditions — and in general there are $m^{k}$ initial conditions for a recurrence of order $k$ — clustering into a variety of irreducible and cyclically equivalent sequences. For example, the period [011011] reduces to [011], which is cyclic equivalent to [101] and [110], capturing the three initial conditions (0,1)(0,1),$(1,1)$ and $(1,0)$.
This naturally leads to a broader question: what is the complete structure of all periodic sequences arising from a given recurrence modulo $m$? How many distinct periods exist and what are their lengths? Understanding this "period landscape" reveals algebraic structure invisible when studying only the canonical Fibonacci sequence.
To approach this systematically, we focus on recurrences with restricted coefficients, namely monic quadratic polynomials with coefficients in $\{-1,+1\}$ (also known as monic Littlewood polynomials [3]). Two of these four polynomials are $x^{2}-x-1$ and $x^{2}+x-1$, which generate the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$. Notably, these are the two deterministic branches of Viswanath’s random recurrence $a_{n}=a_{n-1}\pm a_{n-2}$, where the sign is chosen at random at each step, and where the sequence grows exponentially with rate approximately $1.13198824$ [4]. The other two polynomials are the cyclotomic polynomials $\Phi_{3}(x)=x^{2}+x+1$ and $\Phi_{6}(x)=x^{2}-x+1$, whose roots are primitive roots of unity. More generally, the $n$th cyclotomic polynomial is defined as
$\Phi_{n}(x)={\displaystyle\prod_{\begin{subarray}{c}1\leq k\leq n\\ \gcd(k,n)=1\end{subarray}}\left(x-e^{2i\pi\frac{k}{n}}\right)}$ (3)
satisfying $\prod_{d\mid n}\Phi_{d}(x)=x^{n}-1$. These polynomials are fundamental in algebra and number theory, yet their role in modular periodicity of recurrences remains largely unexplored. Since their roots lie on the unit circle, cyclotomic recurrences exhibit fundamentally different behavior than exponentially growing sequences; they remain bounded, making them a natural starting point for classification.
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences. For the cyclotomic cases, we conjecture explicit formulas for the number of distinct periods modulo $m$ (1,4,7 and 9), verified through extensive computation. Between the Fibonacci and its parity recurrence, we discover an unexpected chiral symmetry: despite having different characteristic roots, they produce identical period structures for all tested moduli. We conjecture that the periodicity of prime moduli is due to the number of zeros in the Pisano period, together with a mysterious variable $\alpha$ (11), and display the fractal-like self-similarity for prime power moduli (16). Furthermore, we conjecture that these periods exhibit weight preservation (18) when extending from divisor moduli to composite moduli, and that both recurrences share an identical minima distribution (21). These findings suggest deep connections between deterministic periodicity in finite rings and questions of exponential growth in random recurrences.
2 Periodicity of Cyclotomic Recurrences
This section addresses the periodicity of recurrences with characteristic polynomial being cyclotomic. For example, following equations 1 and 2, the corresponding recurrence for the first cyclotomic polynomial is $a_{n}=a_{n-1}$, for the third it is $a_{n}=-a_{n-2}-a_{n-1}$, and for the ninth it is $a_{n}=-a_{n-6}-a_{n-3}$. Notably the first cyclotomic polynomial having other coefficients than 1,0, or -1 is $\Phi_{105=3\cdot 5\cdot 7}$.
Conjecture 1 (Period Count for $\Phi_{p}$).
Let $p$ be a prime number and consider the linear recurrence relation of order p1p-1 defined by the characteristic polynomial $\Phi_{p}(x)=\sum_{k=0}^{p-1}x^{k}$, where $\Phi_{p}(x)$ is the $p$-th cyclotomic polynomial. For a given positive integer $m$, let $(\Phi_{p},m)$ denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{p-2})\in(\mathbb{Z}/m\mathbb{Z})^{p-1}$. Then the following formula holds:
$(\Phi_{p},m)=\begin{cases}\frac{m^{p-1}-p}{p}+p&\text{if }p\mid m\\ \frac{m^{p-1}-1}{p}+1&\text{if }p\nmid m\end{cases}$ (4)
Remark 2.
The conjecture implies that when $p$ divides $m$, there are exactly $p$ fixed points (periods of length 1) and $\frac{m^{p-1}-p}{p}$ periods of length $p$. When $p$ does not divide $m$, there is exactly one fixed point (the zero vector) and $\frac{m^{p-1}-1}{p}$ periods of length $p$.
Example 3.
For $p=5$ and $m=10$ (where $5\mid 10$), we have:
$\#_{\pi,5}(10)$ $=\frac{10^{4}-5}{5}+5=\frac{9995}{5}+5=1999+5=2004$ (5)
This corresponds to 5 fixed points (namely [0],[2],[4],[6] and [8]) and 1999 periods of length 5.
Conjecture 4 (Period Count for $\Phi_{2p}$).
Let $p$ be an odd prime number and consider the linear recurrence relation of order p1p-1 defined by the characteristic polynomial $\Phi_{2p}=\sum_{k=0}^{p-1}(-x)^{k}$. For a given positive integer $m$, let $(\Phi_{2p},m)$ denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{n-1})\in(\mathbb{Z}/m\mathbb{Z})^{p-1}$. Then the following formula holds:
$(\Phi_{2p},m)=\begin{cases}\frac{m^{p-1}-p}{2p}+1+\frac{p-1}{2}&\text{if }p\mid m\\ \frac{m^{p-1}-2^{p-1}}{2p}+1+\frac{2^{p-1}-1}{p}&\text{if }2\mid m\\ \frac{m^{p-1}-2^{p-1}-p+1}{2p}+1+\frac{p-1}{2}+\frac{2^{p-1}-1}{p}&\text{if }p\>\text{and }2\mid m\\ \end{cases}$ (6)
Remark 5.
The conjecture implies that when $p$ and 2 divide $m$, there is one fixed point (the zero vector), $\frac{p-1}{2}$ periods of length 2, $\frac{2^{p-1}-1}{p}$ periods of length p and $\frac{m^{p-1}-2^{p-1}-p+1}{2p}$ periods of length $2p$.
Example 6.
For $2p=10$ and $m=10$ (where $p$ and 2 devide m), we have:
$\#_{\pi,10}(10)$ $=\frac{10^{4}-2^{4}-5+1}{10}+1+\frac{4}{2}+\frac{2^{4}-1}{5}=\frac{9980}{10}+6=998+6=1004$ (7)
This corresponds to 1 fixed point, 2 periods of length 2 (namely [28] and [46]), 3 periods of length 5 (namely [00055], [00505] and [05555]) and 998 periods of length 10.
Conjecture 7 (Period Count for $\Phi_{p^{j}}$).
Let $p$ be a prime number and consider the linear recurrence relation of order $p^{j}-p^{j-1}$ defined by the characteristic polynomial $\Phi_{p^{j}}(x)=\sum_{k=0}^{p-1}x^{kp^{j-1}}$, where $\Phi_{p^{j}}(x)$ is the $p^{j}$-th cyclotomic polynomial. For a given positive integer $m$, let $(\Phi_{p^{j}},m)$ denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{p^{j}-p^{j-1}-1})\in(\mathbb{Z}/m\mathbb{Z})^{p^{j}-p^{j-1}}$. Then the following formula holds:
$(\Phi_{p^{j}},m)=\begin{cases}\frac{m^{p^{j}-p^{j-1}}-\sum^{j-1}\limits_{i=0}p^{i}M(p,p^{i})}{p^{j}}+\sum^{j-1}\limits_{i=0}M(p,p^{i})&\text{if }p\mid m\\ \frac{m^{p^{j}-p^{j-1}}-1}{p^{j}}+1&\text{if }p\nmid m\end{cases}$ (8)
, where the periodic lengths are all $p^{i}$ with $i\in\{0,1,2,...,j\}$ if $p\mid m$ and $i\in\{0,j\}$ if $p\nmid m$, and where
$M(m,r)=\frac{1}{r}\sum\limits_{d\mid r}\mu(d)m^{r/d}$ (9)
is the number of different aperiodic $m$-ary necklaces of length $r$, as presented in [5].It is also the number of monic irreducible polynomials of degree r over a finite field $\mathbb{F}_{q}$ (see [6]), with $\mu$ being the classic Möbius function. $M(m,r)$ also refers to Moureau’s necklace-counting function or MacMahon’s formula.
Example 8.
For $p^{j}=3^{2}=9$ and $m=12$ (where $3\mid 12$), we have:
$\#_{\pi,9}(12)$ $=\frac{12^{6}-(1\cdot 3+3\cdot 8)}{9}+(1\cdot 3+3\cdot 8)=\frac{2985957}{9}+27=331773+27=331800$ (10)
This corresponds to 3 fixed points (namely [0],[4] and [8]), 8 periods of length 3 (namely [004], [008], [048],[084], [044], [088], [448] and [884]) and 331773 periods of length 9.
Conjecture 9 (Period Count for $x^{n}-1$).
Consider the linear recurrence relation of order $n$ defined by the characteristic polynomial $x^{n}-1$. For a given positive integer $m$, let #(n,m)\#(n,m) denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{n-1})\in(\mathbb{Z}/m\mathbb{Z})^{n}$. Then the following formula holds:
$(n,m)=\frac{1}{n}\left(m^{n}-\displaystyle\sum_{\begin{subarray}{c}r\mid n\\ r\neq n\end{subarray}}rM(m,r)\right)+\sum_{\begin{subarray}{c}r\mid n\\ r\neq n\end{subarray}}M(m,r)$ (11)
, where the periodic lengths are all $r$ that divide n, and $M(m,r)$ is defined by Equation 9.
Example 10.
For $n=6$ and $m=4$, we have:
$\#^{6}_{\pi}(4)$ $=\frac{4^{6}-(1\cdot 4+2\cdot 6+3\cdot 20)}{6}+(1\cdot 4+2\cdot 6+3\cdot 20)=\frac{4020}{6}+76=670+76=746$ (12)
This corresponds to 4 fixed points (namely [0],[1],[2] and [3]), 6 periods of length 2 (namely [01], [02], [03],[12], [13] and [23]), 20 periods of length 3 (namely [001], [002], [003], [011], [022], [033], [012], [013], [021], [023], [031], [032], [112], [113], [221], [223], [331], [332], [123] and [132]) and 670 periods of length 6.
3 Periodicity of Fibonacci Recurrences
We now turn our focus to the Fibonacci recurrence $a_{n}=a_{n-2}+a_{n-1}$ and its parity transform $a_{n}=a_{n-2}-a_{n-1}$ with random integer initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$. The implied mirror-symmetry can be seen in Table 1 for all moduli $m$, where the number of periods and lengths are similar for both recurrences, and where every period corresponds to a chiral period from the other recurrence, respectively.
Table 1: Periods for different moduli $m$, with a mirror drawn between the two recurrences, and where the Pisano periods are in bold
  m $a_{n}=a_{n-2}+a_{n-1}$ $a_{n}=a_{n-2}-a_{n-1}$
  1 0 0
2 110, 0 0, 011
3 11202210, 0 0, 01220211
4 332130, 112310, 220,0 0, 022, 013211, 031233
5 11230331404432022410, 3421, 0 0, 1243, 01422023440413303211
6 22404420, 330, 0, 112352134150554314532510 0, 033, 02440422, 015235413455051431253211
One open problem in mathematics is calculating the length of the Pisano periods, $\pi(m)$, explicitly. However, classical studies focus on the single canonical sequence, neglecting the complete period landscape arising from all possible initializations. We observe that periods of length 3 emerge at all even moduli, periods of length 8 at every third modulus, and periods of length 4 and 20 at every fifth modulus. These patterns suggest a fundamental principle: the periodicity of composite moduli is completely determined by the periodicity at prime moduli via the Chinese Remainder Theorem. This motivates our systematic classification of prime moduli, which we present below.
Conjecture 11 (Period Count for Prime Fibonacci Recurrences).
Let $p$ be a prime and let #(p)\#(p) denote the number of distinct periods when the Fibonacci recurrence or its parity transform is computed modulo $p$ over all possible initial conditions $(a_{0},a_{1})\in(\mathbb{Z}/p\mathbb{Z})^{2}$. Primes divide into two types:
Type 1 (Two period lengths): The Pisano period satisfies $\pi(p)=\frac{2(p+1)}{\alpha}$ where $\alpha$ is an odd positive integer. The number of distinct periods is:
$(\alpha,p)=\frac{(p^{2}-1)\alpha}{2(p+1)}+1$ (13)
Additionally, for odd and even $\alpha$, the Pisano period satisfies $\pi(p)=\frac{p-1}{\alpha}$ and:
$(\alpha,p)=\frac{(p^{2}-1)\alpha}{p-1}+1$ (14)
Type 2 (Three period lengths): The Pisano period satisfies $\pi(p)=\frac{p-1}{\alpha}$ where $\alpha$ is a positive integer. The number of distinct periods is:
$_{2}(\alpha,p)=p\cdot\alpha+2\alpha+1$ (15)
Type-2 primes are precisely those whose Pisano period contains exactly one zero, corresponding to OEIS sequence A053032 (seperately excluding $p=2$ but including $p=5$).
Remark 12 (Connection to Quadratic Residues).
The classification is intimately connected to the Legendre symbol $\left(\frac{5}{p}\right)$. Type-1 primes with odd $\alpha$ (satisfying $\pi(p)=\frac{2(p+1)}{\alpha}$) correspond precisely to primes where 5 is a quadratic non-residue, while Type-2 primes and Type-1 primes with even $\alpha$ (both satisfying $\pi(p)=\frac{p-1}{\alpha}$) correspond to primes where 5 is a quadratic residue. This follows from the classical result that $\pi(p)\mid(p-1)$ when $\left(\frac{5}{p}\right)=1$ and $\pi(p)\mid 2(p+1)$ when $\left(\frac{5}{p}\right)=-1$.
A finer classification emerges via congruences modulo 20. Primes satisfying $p\equiv 3,7,13,17\pmod{20}$ are Type-1 with $\pi(p)=\frac{2(p+1)}{\alpha}$. Among primes with $\left(\frac{5}{p}\right)=1$, those satisfying $p\equiv 11,19\pmod{20}$ are predominantly Type-2, while $p\equiv 1,9\pmod{20}$ exhibit mixed behavior. The precise determination of $\alpha$ for a given prime remains an open problem.
Remark 13.
For Type-1 primes, the period structure consists of the zero vector (length 1) and #1\#-1 periods of length $\pi(p)$. For Type-2 primes, there are three distinct period lengths: the zero vector (length 1) appears once, length $\frac{\pi(p)}{2}$ appears exactly $2\alpha$ times, and length $\pi(p)$ appears $p\cdot\alpha$ times.
Example 14 (Type-1 Prime: $p=89$, $n=2$).
The Pisano period is $\pi(89)=\frac{89-1}{2}=44$. The number of distinct periods is:
$(2,89)=\frac{(89^{2}-1)\cdot 2}{89-1}+1=\frac{15840}{88}+1=180+1=181$ (16)
The period structure consists of the zero vector and $180$ periods of length $44$.
Example 15 (Type-2 Prime: $p=11$, $n=1$).
The Pisano period is $\pi(11)=\frac{11-1}{1}=10$. The number of distinct periods is:
$(1,11)=11\cdot 1+2\cdot 1+1=14$ (17)
The period structure consists of one period of length $1$ (zero vector), $2$ periods of length $5$, and $11$ periods of length $10$.
Conjecture 16 (Self-Similarity at Prime Powers).
The period structure for the Fibonacci recurrence and its parity transform exhibit hierarchical self-similarity between prime and prime power moduli. At each transition $p^{k}\to p^{k+1}$, all existing periods are preserved, and each period of length $\ell>1$ at $p^{k}$ generates new periods of length $p\ell$ at $p^{k+1}$, with multiplicities scaling by factor $p$. For Type-2 primes, however, the multiplicity for the middle period remains constant at exactly $2\alpha$ across all powers $p^{k}$.
Example 17 (Prime Power Extension: $p=19$, $\alpha=1$).
The base structure at $p=19$ is $\{1,2\times 9,19\times 18\}$ and at $p^{2}=361$ it is $\{1,2\times 9,19\times 18,2\times 171,379\times 342\}$. Observe that the middle period length $9$ scales to $171=9\times 19$ while maintaining count $2\alpha=2$, and that the main period length $\pi(19)=18$ scales to $342=18\times 19$ with adapted multiplicity increasing from $19$ to $379$.
Conjecture 18 (Weight Preservation of Fibonacci Recurrences).
Let $F_{m}$ denote the set of all periodic sequences arising from the Fibonacci recurrence or its parity transform with random initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$, taken modulo $m$. For any divisor $d$ of a composite modulus $m$, we define the weight of a period $p$ of length $\ell_{p}$ in $F_{d}$ as $w_{d}(p)=\frac{\ell_{p}}{d^{2}}$. When a period $p_{d}\in F_{d}$ is extended to $F_{m}$, it gives rise to one or more periods $\{p_{m}^{(1)},\ldots,p_{m}^{(k)}\}\subset F_{m}$ that reduce to $p_{d}$ modulo $d$. We conjecture that the total weight is conserved:
$w_{d}(p_{d})=\sum_{i=1}^{k}w_{m}(p_{m}^{(i)})$ (18)
Example 19 (Extension from $d=2$ to $m=6$).
The space $F_{2}$ contains periods 0 (weight $\frac{1}{4}$) and $011$ (weight $\frac{3}{4}$). When we extend to $F_{6}$, the period 0 extends to periods 0 and $02240442$ with combined weight $\frac{1}{36}+\frac{8}{36}=\frac{9}{36}=\frac{1}{4}$, while the period $011$ extends to periods $033$ and $011235213415055431453251$ with combined weight $\frac{3}{36}+\frac{24}{36}=\frac{27}{36}=\frac{3}{4}$, preserving the original weights exactly.
Example 20 (Extension from $d=3$ to $m=6$).
The space $F_{3}$ contains periods 0 (weight $\frac{1}{9}$) and $01120221$ (weight $\frac{8}{9}$). Upon extension to $F_{6}$, the period 0 extends to periods 0 and $033$ with combined weight $\frac{1}{36}+\frac{3}{36}=\frac{4}{36}=\frac{1}{9}$, while $01120221$ extends to periods $02240442$ and $011235213415055431453251$ with combined weight $\frac{8}{36}+\frac{24}{36}=\frac{32}{36}=\frac{8}{9}$, again preserving the weight distribution perfectly.
Conjecture 21 (Probability distribution of Fibonacci minima).
Consider the Fibonacci recurrence $a_{n}=a_{n-2}+a_{n-1}$ and its chiral recurrence $a_{n}=a_{n-2}-a_{n-1}$ with random integer initialization $\{a_{0},a_{1}\}\in\mathbb{Z}^{2}$. Although these sequences diverge as $n\to\pm\infty$, they possess well-defined absolute minima. Let $F_{n}$ and $L_{n}$ denote the $n$-th Fibonacci and Lucas numbers, respectively. Define $P(n)$ as the probability that a randomly initialized sequence has its absolute minimum at position $n$. Then:
$P(n)=\begin{cases}\frac{1}{4}&\text{if }n=0\\ \frac{1}{\pi}\left(\arctan\left(\frac{L_{n-2}}{L_{n-1}}\right)-\arctan\left(\frac{L_{n}}{L_{n+1}}\right)\right)&\text{if }n>1,\,n\text{ even}\\ \frac{1}{\pi}\left(\arctan\left(\frac{L_{n}}{L_{n+1}}\right)-\arctan\left(\frac{L_{n-2}}{L_{n-1}}\right)\right)&\text{if }n>1,\,n\text{ odd}\end{cases}$ (19)
Moreover, the probability satisfies the symmetry relation P(n)=P(1n)P(n)=P(1-n).
Remark 22.
The probability $P(n)$ arises from the fact, that two minima (of equal absolute value) only exist if $\frac{a_{0}}{a_{1}}=\frac{L_{n}}{L_{n+1}}$ or $\frac{a_{0}}{a_{1}}=\frac{-L_{n}}{L_{n-1}}$ for Fibonaccis recurrence (or if $\frac{a_{0}}{a_{1}}=\frac{L_{n}}{L_{n-1}}$ or $\frac{a_{0}}{a_{1}}=\frac{-L_{n}}{L_{n+1}}$ for its parity transform). Furthermore, between any two consecutive Lucas ratios (e.g. $\frac{L_{n-2}}{L_{n-1}}$ and $\frac{L_{n}}{L_{n+1}}$) lies exactly one mediant Fibonacci ratio (e.g. $\frac{F_{n-1}}{F_{n}}=\frac{L_{n-2}+L_{n}}{L_{n-1}+L_{n+1}}$), which is the unique initial condition for which one $a_{n}=0$ in the recurrence sequence.
Remark 23 (Connection to Modular Forms and Farey Sequences).
The Lucas bounds and intermediate Fibonacci ratios form matrices in $\mathrm{SL}(2,\mathbb{Z})$, suggesting deep connections to modular forms and the theory of continued fractions. The geometric interpretation via arctangent differences may relate to Farey arcs and the tessellation of the upper half-plane by the modular group.
4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect chiral symmetry between the Fibonacci and parity recurrences. We classified primes into two types based on their period structure, observed fractal self-similarity at prime powers, established weight preservation under modular extension, and derived probability distributions for sequence minima governed by Lucas ratios.
A central mystery remains: the classification parameter $\alpha$ that determines a prime’s type and governs the exact period count. We observe that the type classification is intimately connected to whether 5 is a quadratic residue modulo $p$, however, the precise value of $\alpha$ for a given prime remains unknown. Understanding whether $\alpha$ relates to higher-order residue properties, the structure underlying Viswanath’s random recurrence, or other arithmetic invariants could reveal deeper connections between Fibonacci periodicity and algebraic number theory.
The framework extends naturally beyond our quadratic cases. While we investigated the two Littlewood polynomials $x^{2}\pm x-1$ among quadratics, higher-degree polynomials with restricted coefficients also reveal rich periodicity structures. The most significant example are the order-6 recurrences $a_{n}=a_{n-6}+a_{n-3}$ and $a_{n}=a_{n-6}-a_{n-3}$, which are the negations of the cyclotomic recurrences from $\Phi_{9}$ and $\Phi_{18}$ (analogous to how Fibonacci and Parity arise as negations of $\Phi_{3}$ and $\Phi_{6}$). These exhibit identical periodicity for all tested moduli $m\leq 19$, demonstrating that chiral symmetry might extend beyond the quadratic case. Remarkably, classical Fibonacci periods (such as $01120221$ mod 3 and $1342$ and $01123033140443202241$ mod 5) appear within these sequences, suggesting hierarchical connections. However, fully understanding their period structure requires complex computations. Certain primes exhibit five distinct period lengths, rather than the two or three observed in the quadratic case. Additionally, self-similarity at prime powers follows more subtle rules than the direct scaling observed for Fibonacci-Parity. Whether weight preservation and generalized classification formulas extend to this setting, and whether the negations of the cyclotomic recurrences $\Phi_{3^{j}}$ and $\Phi_{2\cdot 3^{j}}$, for all $j\geq 1$, yield chirality remains important open questions. A rigorous understanding may require companion matrix theory and conjugation properties under sign transformations.
Another interesting family is that of the Fekete polynomials, which are related to cyclotomic polynomials and L-functions, as demonstrated in [7]. Studying the complete period landscapes of these and other polynomial families could reveal universal principles governing recurrence periodicity in finite rings.
Finally, preliminary computational evidence suggests intriguing relationships between Fibonacci period lengths and cyclotomic recurrences. Certain Fibonacci periods appear systematically across families of cyclotomic polynomials—for instance, the Pisano period 011 only appears modulo 2 in all $\Phi_{n}$ with $3\mid n$, and modulo 3 in all $\Phi_{3^{j}}$ for all $j>1$. In contrast, periods from primes in OEIS A053032 (Type-2 primes where the Pisano period contains exactly one zero) rarely or never appear in cyclotomic landscapes. Whether this pattern reflects a deeper structural link between exponentially growing sequences (Fibonacci) and unit-root periodicity (cyclotomic) remains an open question worthy of further investigation.