### v2

# Modular Periodicity of Random Initialized Recurrences

Abstract
Classical studies of the Fibonacci sequence focus on its periodicity modulo $m$ (the Pisano periods) with canonical initialization. We investigate instead the complete periodic structure arising from all $m^{2}$ possible initializations in $\mathbb{Z}/m\mathbb{Z}$. We discover perfect chiral symmetry between the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ and observe fractal self-similarity in the extension from prime to prime power moduli. Additionally, we classify prime moduli based on their quadratic reciprocity and demonstrate that periodic sequences exhibit weight preservation under modular extension. Furthermore, we define a minima distribution $P(n)$ governed by Lucas ratios, which satisfies the symmetric relation P(n)=P(1n)P(n)=P(1-n). For cyclotomic recurrences, we propose explicit counting functions for the number of distinct periods with connections to necklace enumeration. These findings imply potential connections to Viswanath’s random recurrence, modular forms and L-functions.
Keywords— Fibonacci sequence, Pisano period, cyclotomic polynomial, chiral symmetry, random recurrence, modular arithmetic
1 Introduction
Linear homogeneous recurrence relations play a fundamental role in number theory and combinatorics. A recurrence of order $k$ is defined as
$a_{n}=r_{1}a_{n-1}+r_{2}a_{n-2}+...+r_{k-1}a_{n-k+1}+r_{k}a_{n-k}=\sum\limits_{i=1}^{k}r_{i}a_{n-i}$ (1)
where the sequence is fully determined by $k$ initial values ($a_{0},a_{1},...,a_{k-1}$) and its characteristic polynomial
$f(x)=x^{k}-r_{1}x^{k-1}-...-r_{k-1}x-r_{k}$ (2)
as shown in [1]. The most celebrated example is the Fibonacci sequence $\{0,1,1,2,3,5,8,13,\ldots\}$, generated by $x^{2}-x-1$ with initial values $a_{0}=0,a_{1}=1$. This sequence grows exponentially, but when reduced modulo $m$, it becomes periodic with period length $\pi(m)$, known as the Pisano period [2]. Determining $\pi(m)$ explicitly remains an open problem in number theory.
However, the classical Pisano problem addresses only a single sequence with fixed initialization. In $\mathbb{Z}/m\mathbb{Z}$, there are $m^{2}$ possible initial conditions — and in general there are $m^{k}$ initial conditions for a recurrence of order $k$ — clustering into a variety of irreducible and cyclically equivalent sequences. For example, the period [011011] reduces to [011], which is cyclic equivalent to [101] and [110], capturing the three initial conditions (0,1)(0,1),$(1,1)$ and $(1,0)$. This naturally leads to a broader question: what is the complete structure of all periodic sequences arising from a given recurrence modulo $m$? How many distinct periods exist and what are their lengths? Understanding this "period landscape" reveals algebraic structure invisible when studying only the canonical Fibonacci sequence.
To approach this systematically, we focus on recurrences with restricted coefficients, namely monic quadratic polynomials with coefficients in $\{-1,+1\}$ (also known as Littlewood polynomials [3]). Two of these four polynomials are $x^{2}-x-1$ and $x^{2}+x-1$, which generate the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$. The other two are the cyclotomic polynomials $\Phi_{3}(x)=x^{2}+x+1$ and $\Phi_{6}(x)=x^{2}-x+1$, whose roots are primitive roots of unity. Interestingly, the Fibonacci recurrence (which is the negation from $\Phi_{3}$) and the recurrence $a_{n}=a_{n-1}-a_{n-2}$ (from $\Phi_{6}$) form the two deterministic branches of Viswanath’s random recurrence $a_{n}=a_{n-1}\pm a_{n-2}$, which exhibits exponential growth with rate approximately $1.13198824$ [4]. This connection between non-cyclotomic and cyclotomic recurrences motivates our parallel investigation of both families. More generally, the $n$th cyclotomic polynomial is defined as
$\Phi_{n}(x)={\displaystyle\prod_{\begin{subarray}{c}1\leq k\leq n\\ \gcd(k,n)=1\end{subarray}}\left(x-e^{2i\pi\frac{k}{n}}\right)}$ (3)
satisfying $\prod_{d\mid n}\Phi_{d}(x)=x^{n}-1$. These polynomials are fundamental in algebra and number theory, yet their role in modular periodicity of recurrences remains largely unexplored. Since their roots lie on the unit circle, cyclotomic recurrences exhibit fundamentally different behavior than exponentially growing sequences; they remain bounded, making them a natural starting point for classification.
This paper investigates the period landscapes of both cyclotomic and non-cyclotomic recurrences across three main sections. In section 2, we conjecture explicit counting functions for the number of distinct periods modulo $m$ for various cyclotomic families, including $\Phi_{p}$ (1), $\Phi_{2p}$ (4), $\Phi_{p^{j}}$ (7) and $x^{n}-1$ (9), all supported by extensive computational verification. In section 3, the non-cyclotomic Fibonacci and parity recurrences are explored, revealing unexpected chiral symmetry. Despite having algebraically distinct characteristic roots, these recurrences produce identical period structures for all tested moduli. We classify prime moduli based on their quadratic reciprocity and an additional parameter $\alpha$ (11), observe fractal self-similarity at prime power moduli (15), and conjecture weight preservation under modular extension (17). Furthermore, we establish that both recurrences share an identical minimum distribution governed by Lucas number ratios (20). In section 4, we discuss broader implications and open questions, and suggest that deterministic periodicity in finite rings has potential relations to stochastic growth in random recurrences and to classical number-theoretical objects such as modular forms and L-functions.
2 Periodicity of Cyclotomic Recurrences
This section addresses the periodicity of recurrences with characteristic polynomial being cyclotomic. For example, following equations 1 and 2, the corresponding recurrence for the first cyclotomic polynomial is $a_{n}=a_{n-1}$, for the third it is $a_{n}=-a_{n-1}-a_{n-2}$ and for the ninth it is $a_{n}=-a_{n-3}-a_{n-6}$. Notably the first cyclotomic polynomial having other coefficients than 1,0, or -1 is $\Phi_{105=3\cdot 5\cdot 7}$.
Conjecture 1 (Period Count for $\Phi_{p}$).
Let $p$ be a prime number and consider the linear recurrence relation of order p1p-1 defined by the characteristic polynomial $\Phi_{p}(x)=\sum_{k=0}^{p-1}x^{k}$, where $\Phi_{p}(x)$ is the $p$-th cyclotomic polynomial. For a given positive integer $m$, let $(\Phi_{p},m)$ denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{p-2})\in(\mathbb{Z}/m\mathbb{Z})^{p-1}$. Then the following formula holds:
$(\Phi_{p},m)=\begin{cases}\frac{m^{p-1}-p}{p}+p&\text{if }p\mid m\\ \frac{m^{p-1}-1}{p}+1&\text{if }p\nmid m\end{cases}$ (4)
Remark 2.
The conjecture implies that when $p$ divides $m$, there are exactly $p$ fixed points (periods of length 1) and $\frac{m^{p-1}-p}{p}$ periods of length $p$. When $p$ does not divide $m$, there is exactly one fixed point (the zero vector) and $\frac{m^{p-1}-1}{p}$ periods of length $p$.
Example 3.
For $p=5$ and $m=10$ (where $5\mid 10$), we have:
$\#(\Phi_{5},10)$ $=\frac{10^{4}-5}{5}+5=\frac{9995}{5}+5=1999+5=2004$ (5)
This corresponds to 5 fixed points (namely [0],[2],[4],[6] and [8]) and 1999 periods of length 5.
Conjecture 4 (Period Count for $\Phi_{2p}$).
Let $p$ be an odd prime number and consider the linear recurrence relation of order p1p-1 defined by the characteristic polynomial $\Phi_{2p}=\sum_{k=0}^{p-1}(-x)^{k}$. For a given positive integer $m$, let $(\Phi_{2p},m)$ denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{n-1})\in(\mathbb{Z}/m\mathbb{Z})^{p-1}$. Then the following formula holds:
$(\Phi_{2p},m)=\begin{cases}\frac{m^{p-1}-p}{2p}+1+\frac{p-1}{2}&\text{if }p\mid m\\ \frac{m^{p-1}-2^{p-1}}{2p}+1+\frac{2^{p-1}-1}{p}&\text{if }2\mid m\\ \frac{m^{p-1}-2^{p-1}-p+1}{2p}+1+\frac{p-1}{2}+\frac{2^{p-1}-1}{p}&\text{if }p\>\text{and }2\mid m\\ \end{cases}$ (6)
Remark 5.
The conjecture implies that when $p$ and 2 divide $m$, there is one fixed point (the zero vector), $\frac{p-1}{2}$ periods of length 2, $\frac{2^{p-1}-1}{p}$ periods of length p and $\frac{m^{p-1}-2^{p-1}-p+1}{2p}$ periods of length $2p$.
Example 6.
For $2p=10$ and $m=10$ (where $p$ and 2 devide m), we have:
$\#(\Phi_{10},10)$ $=\frac{10^{4}-2^{4}-5+1}{10}+1+\frac{4}{2}+\frac{2^{4}-1}{5}=\frac{9980}{10}+6=998+6=1004$ (7)
This corresponds to 1 fixed point, 2 periods of length 2 (namely [28] and [46]), 3 periods of length 5 (namely [00055], [00505] and [05555]) and 998 periods of length 10.
Conjecture 7 (Period Count for $\Phi_{p^{j}}$).
Let $p$ be a prime number and consider the linear recurrence relation of order $p^{j}-p^{j-1}$ defined by the characteristic polynomial $\Phi_{p^{j}}(x)=\sum_{k=0}^{p-1}x^{kp^{j-1}}$, where $\Phi_{p^{j}}(x)$ is the $p^{j}$-th cyclotomic polynomial. For a given positive integer $m$, let $(\Phi_{p^{j}},m)$ denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{p^{j}-p^{j-1}-1})\in(\mathbb{Z}/m\mathbb{Z})^{p^{j}-p^{j-1}}$. Then the following formula holds:
$(\Phi_{p^{j}},m)=\begin{cases}\frac{m^{p^{j}-p^{j-1}}-\sum^{j-1}\limits_{i=0}p^{i}M(p,p^{i})}{p^{j}}+\sum^{j-1}\limits_{i=0}M(p,p^{i})&\text{if }p\mid m\\ \frac{m^{p^{j}-p^{j-1}}-1}{p^{j}}+1&\text{if }p\nmid m\end{cases}$ (8)
, where the periodic lengths are all $p^{i}$ with $i\in\{0,1,2,...,j\}$ if $p\mid m$ and $i\in\{0,j\}$ if $p\nmid m$, and where
$M(m,r)=\frac{1}{r}\sum\limits_{d\mid r}\mu(d)m^{r/d}$ (9)
is the number of different aperiodic $m$-ary necklaces of length $r$, as presented in [5].It is also the number of monic irreducible polynomials of degree r over a finite field $\mathbb{F}_{q}$ (see [6]), with $\mu$ being the classic Möbius function. $M(m,r)$ also refers to Moureau’s necklace-counting function or MacMahon’s formula.
Example 8.
For $p^{j}=3^{2}=9$ and $m=12$ (where $3\mid 12$), we have:
$\#(\Phi_{9},12)$ $=\frac{12^{6}-(1\cdot 3+3\cdot 8)}{9}+(3+8)=\frac{2985957}{9}+11=331773+27=331784$ (10)
This corresponds to 3 fixed points (namely [0],[4] and [8]), 8 periods of length 3 (namely [004], [008], [044],[088], [048], [084], [448] and [884]) and 331773 periods of length 9.
Conjecture 9 (Period Count for $x^{n}-1$).
Consider the linear recurrence relation of order $n$ defined by the characteristic polynomial $x^{n}-1$. For a given positive integer $m$, let #(n,m)\#(n,m) denote the number of distinct periods when the recurrence is computed modulo $m$ over all possible initial conditions $(a_{0},a_{1},\ldots,a_{n-1})\in(\mathbb{Z}/m\mathbb{Z})^{n}$. Then the following formula holds:
$(n,m)=\frac{1}{n}\left(m^{n}-\displaystyle\sum_{\begin{subarray}{c}r\mid n\\ r\neq n\end{subarray}}rM(m,r)\right)+\sum_{\begin{subarray}{c}r\mid n\\ r\neq n\end{subarray}}M(m,r)$ (11)
, where the periodic lengths are all $r$ that divide n, and $M(m,r)$ is defined by Equation 9.
Example 10.
For $n=6$ and $m=4$, we have:
$\#(6,4)$ $=\frac{4^{6}-(1\cdot 4+2\cdot 6+3\cdot 20)}{6}+(4+6+20)=\frac{4020}{6}+30=670+30=700$ (12)
This corresponds to 4 fixed points (namely [0],[1],[2] and [3]), 6 periods of length 2 (namely [01], [02], [03],[12], [13] and [23]), 20 periods of length 3 (namely [001], [002], [003], [011], [022], [033], [012], [013], [021], [023], [031], [032], [112], [113], [221], [223], [331], [332], [123] and [132]) and 670 periods of length 6.
3 Periodicity of Fibonacci Recurrences
We now turn our focus to the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its parity transform $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$. The implied mirror-symmetry can be seen in Table 1 for all moduli $m$, where the number of periods and lengths are similar for both recurrences, and where every period corresponds to a chiral period from the other recurrence.
Table 1: Periods for different moduli $m$, with a mirror drawn between the two recurrences, and where the Pisano periods are in bold
  m $a_{n}=a_{n-1}+a_{n-2}$ $a_{n}=-a_{n-1}+a_{n-2}$
  1 0 0
2 110, 0 0, 011
3 11202210, 0 0, 01220211
4 332130, 112310, 220,0 0, 022, 013211, 031233
5 11230331404432022410, 3421, 0 0, 1243, 01422023440413303211
6 22404420, 330, 0, 112352134150554314532510 0, 033, 02440422, 015235413455051431253211
One open problem in mathematics is calculating the length of the Pisano periods, $\pi(m)$, explicitly. However, classical studies focus on the single canonical sequence, neglecting the complete period landscape arising from all possible initializations. We observe that periods of length 3 emerge at all even moduli, periods of length 8 at every third modulus, and periods of length 4 and 20 at every fifth modulus. These patterns suggest a fundamental principle: the periodicity of composite moduli is completely determined by the periodicity at prime moduli via the Chinese Remainder Theorem. This motivates our systematic classification of prime moduli, which we present below.
Conjecture 11 (Period Count for Fibonacci Recurrences modulo p).
Let $p$ be a prime and let #(p)\#(p) denote the number of distinct periods when the Fibonacci recurrence or its parity transform is computed modulo $p$ over all possible initial conditions $(a_{0},a_{1})\in(\mathbb{Z}/p\mathbb{Z})^{2}$. The count is determined by the Legendre symbol $\left(\frac{5}{p}\right)$ and a positive integer $\alpha$ governing the Pisano period:
[…]
$_{A}(p)=\frac{\alpha}{2}(p-1)+1$ (13)
These primes exhibit two period lengths: the zero vector and non-trivial periods of length $\pi_{A}(p)$.
[…]
Subclass B1 (two period lengths): The number of distinct periods is:
$_{B1}(p)=\alpha(p+1)+1$ (14)
These primes exhibit two periodic lengths: the zero vector and non-trivial periods of length $\pi_{B}(p)$.
Subclass B2 (three period lengths): The Pisano period contains exactly one zero (OEIS A053032 for $p\geq 11$). The number of distinct periods is:
$_{B2}(p)=\alpha(p+2)+1$ (15)
These primes exhibit three period lengths: the zero vector, an intermediate length $\frac{\pi_{B}(p)}{2}$ appearing $2\alpha$ times, and the length $\pi_{B}(p)$ appearing $p\alpha$ times.
All primes satisfying $p\equiv 11,19\pmod{20}$ belong to subclass B2, while those satisfying $p\equiv 1,9\pmod{20}$ can belong to either subclass.
Example 12 (Class A Prime: $p=47$, $\alpha=3$).
The Pisano period is $\pi_{A}(47)=\frac{2(47+1)}{3}=32$. The number of distinct periods is:
$_{A}(47)=\frac{3}{2}(47-1)+1=69+1=70$ (16)
The period structure consists of the zero vector and $69$ periods of length $32$.
Example 13 (Class B1 Prime: $p=89$, $\alpha=2$).
The Pisano period is $\pi_{B}(89)=\frac{89-1}{2}=44$. The number of distinct periods is:
$_{B1}(89)=2\cdot(89+1)+1=180+1=181$ (17)
The period structure consists of the zero vector and $180$ periods of length $44$.
Example 14 (Class B2 Prime: $p=11$, $\alpha=1$).
The Pisano period is $\pi_{B}(11)=\frac{11-1}{1}=10$. The number of distinct periods is:
$_{B2}(11)=1\cdot(11+2)+1=13+1=14$ (18)
The period structure consists of the zero vector, $2$ periods of length $5$ and $11$ periods of length $10$.
Conjecture 15 (Self-Similarity at Prime Powers).
The period structure for the Fibonacci recurrence and its parity transform exhibit hierarchical self-similarity between prime and prime power moduli. At each transition $p^{k}\to p^{k+1}$, all existing periods are preserved and each period of length $\ell>1$ at $p^{k}$ generates new periods of length $p\ell$ at $p^{k+1}$, with multiplicities scaling by factor $p$. For class B2 primes, however, the multiplicity for the middle period remains constant at exactly $2\alpha$ across all powers $p^{k}$.
Example 16 (Prime Power Extension for B2 prime $p=19$, $\alpha=1$).
The base structure at $p=19$ is $\{1,2\times 9,19\times 18\}$ and at $p^{2}=361$ it is $\{1,2\times 9,19\times 18,2\times 171,379\times 342\}$. Observe that the middle period length $9$ scales to $171=9\times 19$ while maintaining count $2\alpha=2$, and that the main period length $\pi(19)=18$ scales to $342=18\times 19$ with adapted multiplicity increasing from $19$ to $379$.
Conjecture 17 (Weight Preservation of Fibonacci Recurrences).
Let $F_{m}$ denote the set of all periodic sequences arising from the Fibonacci recurrence or its parity transform with random initialization $(a_{0},a_{1})\in(\mathbb{Z}/m\mathbb{Z})^{2}$, taken modulo $m$. For any divisor $d$ of a composite modulus $m$, we define the weight of a period $p$ of length $\ell_{p}$ in $F_{d}$ as $w_{d}(p)=\frac{\ell_{p}}{d^{2}}$. When a period $p_{d}\in F_{d}$ is extended to $F_{m}$, it gives rise to one or more periods $\{p_{m}^{(1)},\ldots,p_{m}^{(k)}\}\subset F_{m}$ that reduce to $p_{d}$ modulo $d$. We conjecture that the total weight is conserved:
$w_{d}(p_{d})=\sum_{i=1}^{k}w_{m}(p_{m}^{(i)})$ (19)
Example 18 (Extension from $d=2$ to $m=6$).
The space $F_{2}$ contains periods 0 (weight $\frac{1}{4}$) and $011$ (weight $\frac{3}{4}$). When we extend to $F_{6}$, the period 0 extends to periods 0 and $02240442$ with combined weight $\frac{1}{36}+\frac{8}{36}=\frac{9}{36}=\frac{1}{4}$, while the period $011$ extends to periods $033$ and $011235213415055431453251$ with combined weight $\frac{3}{36}+\frac{24}{36}=\frac{27}{36}=\frac{3}{4}$, preserving the original weights exactly.
Example 19 (Extension from $d=3$ to $m=6$).
The space $F_{3}$ contains periods 0 (weight $\frac{1}{9}$) and $01120221$ (weight $\frac{8}{9}$). Upon extension to $F_{6}$, the period 0 extends to periods 0 and $033$ with combined weight $\frac{1}{36}+\frac{3}{36}=\frac{4}{36}=\frac{1}{9}$, while $01120221$ extends to periods $02240442$ and $011235213415055431453251$ with combined weight $\frac{8}{36}+\frac{24}{36}=\frac{32}{36}=\frac{8}{9}$, again preserving the weight distribution perfectly.
Conjecture 20 (Probability distribution of Fibonacci minima).
Consider the Fibonacci recurrence $a_{n}=a_{n-1}+a_{n-2}$ and its chiral recurrence $a_{n}=-a_{n-1}+a_{n-2}$ with random integer initialization $\{a_{0},a_{1}\}\in\mathbb{Z}^{2}$. Although these sequences diverge as $n\to\pm\infty$, they possess well-defined absolute minima. Let $F_{n}$ and $L_{n}$ denote the $n$-th Fibonacci and Lucas numbers respectively. Define $P(n)$ as the probability that a randomly initialized sequence has its absolute minimum at position $n$. Then:
$P(n)=\begin{cases}\frac{1}{4}&\text{if }n=0\\ \frac{1}{\pi}\left(\arctan\left(\frac{L_{n-2}}{L_{n-1}}\right)-\arctan\left(\frac{L_{n}}{L_{n+1}}\right)\right)&\text{if }n>1,\,n\text{ even}\\ \frac{1}{\pi}\left(\arctan\left(\frac{L_{n}}{L_{n+1}}\right)-\arctan\left(\frac{L_{n-2}}{L_{n-1}}\right)\right)&\text{if }n>1,\,n\text{ odd}\end{cases}$ (20)
Moreover, the probability satisfies the symmetry relation P(n)=P(1n)P(n)=P(1-n).
Remark 21.
The probability $P(n)$ arises from the fact, that two minima (of equal absolute value) only exist if $\frac{a_{0}}{a_{1}}=\frac{L_{n}}{L_{n+1}}$ or $\frac{a_{0}}{a_{1}}=\frac{-L_{n}}{L_{n-1}}$ for Fibonaccis recurrence (or if $\frac{a_{0}}{a_{1}}=\frac{L_{n}}{L_{n-1}}$ or $\frac{a_{0}}{a_{1}}=\frac{-L_{n}}{L_{n+1}}$ for its parity transform). Furthermore, between any two consecutive Lucas ratios (e.g. $\frac{L_{n-2}}{L_{n-1}}$ and $\frac{L_{n}}{L_{n+1}}$) lies exactly one mediant Fibonacci ratio (e.g. $\frac{F_{n-1}}{F_{n}}=\frac{L_{n-2}+L_{n}}{L_{n-1}+L_{n+1}}$), which is the unique initial condition for which one $a_{n}=0$ in the recurrence sequence.
Remark 22 (Connection to Modular Forms and Farey Sequences).
The Lucas bounds and intermediate Fibonacci ratios form matrices in $\mathrm{SL}(2,\mathbb{Z})$, suggesting deep connections to modular forms and the theory of continued fractions. The geometric interpretation via arctangent differences may relate to Farey arcs and the tessellation of the upper half-plane by the modular group.
4 Perspectives
This paper examined the modular periodicity of randomly initialized Fibonacci and cyclotomic recurrences, revealing previously unknown symmetries. We conjectured explicit counting formulas for cyclotomic recurrences and discovered perfect chiral symmetry between the Fibonacci and parity recurrences. We classified prime moduli based on their quadratic reciprocity, observed fractal self-similarity at prime power moduli, established weight preservation under modular extension and derived probability distributions for sequence minima governed by Lucas ratios.
The framework of random initialization naturally extends beyond our quadratic cases. While we investigated the two Littlewood polynomials $x^{2}\pm x-1$ among quadratics, higher-degree polynomials with restricted coefficients also reveal rich periodicity structures. The most significant examples are the order-6 recurrences $a_{n}=a_{n-3}+a_{n-6}$ and $a_{n}=-a_{n-3}+a_{n-6}$ (which arise as negations from $\Phi_{9}$ and $\Phi_{18}$). Analogous to how the Fibonacci and parity recurrences are negations from $\Phi_{3}$ and $\Phi_{6}$, these order-6 recurrences also exhibit identical periodicity for all tested moduli $m\leq 19$, demonstrating that chiral symmetry might extend beyond the quadratic case. A rigorous proof of such symmetries may require companion matrix theory and conjugation properties under sign transformations. Observations show that classical Fibonacci periods (such as $01120221$ mod 3 and $1342$ and $01123033140443202241$ mod 5) appear within these order-6 recurrences, suggesting hierarchical connections. However, their period structure is more complex: certain primes exhibit five distinct period lengths rather than the two or three observed in the quadratic case, and self-similarity at prime powers follows subtler rules. Whether weight preservation and generalized classification formulas extend to this setting remains an open question. Moreover, both order-6 recurrences can be viewed as three interleaved Fibonacci or parity recurrences respectively (via initial ratios $a_{0}/a_{3}$, $a_{1}/a_{4}$ and $a_{2}/a_{5}$), suggesting that the Lucas ratio framework and connections to modular forms established for the quadratic case may generalize, though deriving explicit minimum distributions for higher-order systems remains unexplored.
Beyond these generalizations, the interplay between different recurrence families appears to encode deep structural information. Preliminary computation reveals systematic patterns in how Fibonacci periods distribute across cyclotomic recurrences: for instance, the period $011$ appears modulo 2 in all $\Phi_{n}$ with $3\mid n$, and modulo 3 in all $\Phi_{3^{j}}$ for $j\geq 2$, while periods from class B2 primes (OEIS A053032, $p\geq 11$) rarely or never appear in cyclotomic landscapes. Whether this reflects deeper structural incompatibility between exponentially growing sequences and unit-root periodicity remains open. Studying complete period landscapes across polynomial families could reveal universal principles governing recurrence periodicity in finite rings. Fekete polynomials, which relate to cyclotomic polynomials and L-functions, as shown in [7], represent a particularly promising candidate for such investigations.
Finally, a central mystery remains: the classification parameter $\alpha$ that determines a prime’s class and governs the exact period count. While we established that the classification is intimately connected to the Legendre symbol $\left(\frac{5}{p}\right)$, with finer structure emerging via congruences modulo 20, the precise value of $\alpha$ for a given prime remains unknown. Understanding whether $\alpha$ relates to higher-order residue properties, Viswanath’s random recurrence through spectral properties, or other arithmetic invariants could eventually lead to an explicit description of the Pisano periods—resolving one of the longstanding open problems in the theory of Fibonacci sequences.